Find the real number x and y such that 3x+2iy-1x+5y=7+5i
1
Expert's answer
2019-10-18T09:36:51-0400

"3x+2iy-x+5y=7+5i,"

"(2x+5y)+2iy=7+5i,"

"2x+5y=7,\;\;2y=5,"

"y=\frac{5}{2},\;\;x=\frac{1}{2}(7-5*\frac{5}{2})=-\frac{11}{4}."

So, "x=-\frac{11}{4},\;\;y=\frac{5}{2}" .